OpenStudy (anonymous):

What is the slope of a line that is perpendicular to 6x + 3y = –9? A. 2 B. 1/2 C. –2 D. -1/2

OpenStudy (lucaz):

divide the entire equation by three and isolate y y=-2x-3, this is the slope intercept form y=m(x)+b where m=slope and b=y-intercept slopes of perpendicular lines are the negative reciprocal of each other here m = -2, the neg. rec. is -1/m = -1/-2 = 1/2

OpenStudy (anonymous):

so its b ?

OpenStudy (lucaz):

yes, slope is 1/2